OpenStudy (anonymous):

The measure of one angle is 47(degrees) less than the measure of its supplement. Find the measure of each angle.

OpenStudy (mathstudent55):

What are supplementary angles?

OpenStudy (anonymous):

180 degrees

OpenStudy (anonymous):

x=(180-x)-47

OpenStudy (mathstudent55):

Right. Supplementary angles are angles whose measures add up to 180 degrees. If one angle measures x and its supplement measures y, then x + y = 180. Ok?

OpenStudy (anonymous):

yeah i know that i just didnt get dis question set up

OpenStudy (mathstudent55):

Now try to express the angle of measure y in terms of x. Start with x + y = 180 and solve for y: x + y = 180 -x - x ------------------ y = 180 - x Now you see that for two supplementary angles, if one is x, the other one is 180 -x.

OpenStudy (mathstudent55):

Now you follow what the problem tells you: The measure of one angle is 47(degrees) less than the measure of its supplement. x = 180 - x - 47 or simply: x = 180 - x - 47 Solve for x. then find 180 - x to find the measure of the supplement.

OpenStudy (anonymous):

so i got to 2x=133

OpenStudy (mathstudent55):

Good. One more step to get x.

OpenStudy (anonymous):

Oh ok i got it thx man

OpenStudy (mathstudent55):

wlcm